A electric model train travels at 34.7 cm/s around a circular track of radius 1.67 m. How many revolutions does it perform per second, i.e, what is the motion\'s frequency?

Find the train\'s period of revolution.
Answer #1

Distance for complete revolution= perimeter of track

                                    =2*pi*r=2*3.14*1.67=10.493 m

Speed of train=34.7 cm/s=0.347 m/s

Revolution performed per sec=speed/ Distance in one rev =10.493/0.347

                                    =0.033 rev/sec

Period of rev= 1/0.033=30.24 sec